# SReferences

### v2

@@ 5.1 Overview and basic data analysis
We apply the framework developed in this paper to analyze the market-wide extremal dependence of China’s futures market. The data are represented as a multivariate time series of percentage price changes of futures for commodities with high liquidity. The return is calculated at 30-second intervals and the detailed procedure of data generation and access are given in the Supplementary Material, Section B. We consider a subset of the dataset containing 55 assets, which covers a wide range of commodities, including agricultural products, base metals, and ferrous metals. We select a one-year period of data from 2022-08-01 to 2023-07-31, which contains $T=\;$$110,985$ observations.
For each of the $55$ assets, we study bulk and tail temporal dependence by estimating their autocorrelation and extremogram (Davis et al.,, 2012) functions, respectively. The estimated (extremal) dependence decays rapidly, dropping below values of $0.05$ and $0.1$ (for the auto-correlation and extremogram, respectively) after $10$ time lags. Therefore, to remove temporal dependence from the data, we subsample every tenth observation. A detailed description of this dependence analysis is deferred to the Supplementary Material, Section D.1. After sub-sampling, we transform the data to be balanced regularly varying. In the Supplementary Material, Section D.2, we investigate the marginal tail behaviour of the data and find reasonable evidence to suggest that the upper and lower tails of each asset are regularly varying; we thus employ the first marginal transformation as described in Section 3.2. Unreported experiments found this choice to have negligible effect on the final results. We denote the transformed data by $\bm{Z}_{t}=(Z_{1,t},\dots,Z_{55,t})^{\top}$ for $t=1,\dots,T$.
The rest of this section is organized as follows: In Section 5.2, we visualize extremal dependence in the market using the extremal ball. In Section 5.3, we test the ETH for China’s futures market and interpret our findings. In Section 5.4, we capitalize on our test results to construct an artificial dynamic portfolio consisting of tail-inefficient assets and backtest an investment strategy on out-of-sample data. We also study the uncertainty arising from the marginal transformation via a bootstrap experiment, which is deferred to Supplementary Material, Section D.4.
@@ 5.2 Visualization of market-wide extremal dependence
We observe that most of the points concentrate on the tail efficient line, covering a wide range of strengths of extremal dependence. This suggests that most one-step forward extremal dependencies are directionally tail symmetric. Although some extreme events appear to be closely linked to subsequent large movements in asset prices, this connection does not necessarily yield a clear directional signal. At best, it serves as an indicator of heightened volatility, implying that the market remains relatively efficient even under extreme conditions. A similar visualization is conducted for the contemporaneous extremal dependence structure which reveals the pattern of concurrence of extremal events across assets; we defer this to Supplementary Material, Section D.3 for brevity.
Figure 4: Extremal ball for one-step forward extremal dependencies in China’s futures market. Significant tail-inefficient pairs with p-values smaller than 0.01 are circled (based on the permutation test with $\hat{\lambda}^{2}_{n}$). Red points denote pairs $(Z_{i,t},Z_{j,t+1})^{\top}$ and blue points denote $(-Z_{i,t},Z_{j,t+1})^{\top}$.
5.3 Testing the Efficient Tail Hypothesis for China’s futures market
@@
By analyzing the sign of the explanatory variable $X_{i}$ and the sign of $\lambda(X_{i},Y_{j})$, we can classify the significant pairs into two types: momentum effects, where ${\rm Sgn}(X_{i})={\rm Sgn}(\lambda(X_{i},Y_{j}))$, and contrarian effects, otherwise. Momentum (contrarian) effects are well-known phenomena in financial markets, reflecting a tendency for price changes to follow the same (different) directional price movements; see e.g., Asness et al., (2013) for more details. The phenomenon of momentum in extremes may suggest that the market is under-active in extreme situations and indicates the potential contagion path of extreme risk when the market undergoes stress. The phenomenon of contrarian in extremes may suggest that the market is over-active during extreme events and tends to recover in the following periods.
We also analyze the in-product (i.e., the same asset for $X_{i}$ and $Y_{j}$ ) and cross-product (i.e., different assets) pairs. Table 1 shows that the majority of the significant pairs are cross-product pairs (8 out of 10).
Statistics on all 60 significant pairs exhibit similar patterns: 41 pairs are driven by negative extreme events and 19 pairs by positive extreme events; 41 pairs are driven by momentum effects and 19 by contrarian effects. There are only three in-product significant pairs: egg futures (driven by an extremal loss), natural rubber (driven by an extremal gain), and stainless steel (driven by an extremal gain). These three pairs are all contrarian in extremes, while the remaining 57 pairs are cross-product pairs. Within the cross-product pairs, 41 pairs are driven by momentum effects.
Overall, the findings suggest that negative extreme movements are more influential than positive ones, and that momentum effects in extremes tend to appear in cross-product pairs whereas contrarian effects tend to appear in in-product pairs.
Such statistics show that, in the context of extremes of China’s futures market, an asset’s own price often exhibits an initial overreaction, followed by a subsequent correction, which may form the basis of contrarian effects in extremes for tail-inefficient in-product pairs. As for the momentum effects in extremes for the cross-product pairs, as mentioned in Hou, (2007), the speed at which the price of an asset reacts to new information varies among products. The “slow information diffusion hypothesis”, as described by Hou, (2007), may explain our observed momentum effects in the cross-product pairs. One plausible explanation is that certain assets swiftly incorporate relevant news, leading to initial price adjustments in explanatory assets, while others reflect the same information only after a delay. However, it remains an open question as to whether the observed cross-product linkages in extremes reflect truly causal relationships.
@@ 5.4 Exploiting the Efficient Tail Hypothesis analysis
For each pair satisfying $|\hat{\lambda}_{n}^{2}(X_{i},Y_{j})|>0$, we suppose that we enter into the market for asset $Y_{j}$ (long position if $\hat{\lambda}_{n}^{2}(X_{i},Y_{j})>0$, otherwise short position) when we observe an extreme value for asset $X_{i}$ (given it exceeds its 99.5% historical quantile for positive extremes and falls below its 0.5% quantile when considering negative extremes). We hold the asset for one period and then exit the market by closing the position. With the trading cost ignored, we backtest this strategy on the out-of-sample data from 2023-06-01 to 2023-07-31. The cumulative Profit and Loss for each of the 60 significant pairs is given in Figure 5, alongside the overall Profit and Loss for the portfolio. We also study the case when set significance level to $0.05$ under which we find 294 significant pairs. Both results show a positive Profit and Loss throughout, suggesting that a trading strategy based on tail inefficiencies has the potential to make excess returns. This result is consistent with the bootstrap experiment, which is deferred to Supplementary Material, Section D.4.
As stated by the Efficient Market Hypothesis, past information cannot be used to predict future price movements. However, our backtesting on out-of-sample data suggests that our trading strategy using tail-inefficient assets is profitable (under the ideal situation where the trading cost is not considered), which gives evidence against the Efficient Market Hypothesis. Therefore, the Efficient Tail Hypothesis, which is testable, can be viewed as an analogue of the Efficient Market Hypothesis with a stronger focus on extremal events.
Figure 5: Profit and Loss of a dynamic portfolio consisting of the 60 or 294 tail-inefficient pairs under significance level $\alpha^{*}=0.01$ (left) or $\alpha^{*}=0.05$ (right), respectively. Upper panels show the Profit and Loss for each pair and the lower panels show the running cumulative Profit and Loss for the portfolio. The $y$-axis is the cumulative Profit and Loss (%) and the $x$-axis is the time index for the two-month test period.
@@ 6 Conclusion
In this paper, we studied the pairwise directional tail dependence asymmetry for components of a random vector, $\bm{X}$. Under balanced regular variation of $\bm{X}$ over the entire $\mathbb{R}^{p}$, we study the theoretical properties characterizing the dependence structure of $\bm{X}$ in all lower/upper tail quadrants, and we then create a novel tool, the extremal ball, to visualize the pairwise directional asymmetry of tail dependence within high-dimensional random vectors. Furthermore, we proved constraints on the tail dependence in two adjacent quadrants. Such a theoretical result suggests that knowledge of extremal dependence can aid in understanding the extremal dependence in adjacent quadrants of $\mathbb{R}^{p}$.
@@
We proposed a new measure, the directional tail dependence $\lambda(X,Y)$, to quantify the extent of directional tail symmetry. Two estimators are proposed for $\lambda(X,Y)$. We proved asymptotic normality for the first estimator; deriving the asymptotic properties for the second estimator remains future work. We further developed a permutation test for the null hypothesis $\mathcal{H}_{0}:\lambda(X,Y)=0$ and verified its effectiveness and power through simulation studies. We showcased that the permutation test with the second estimator $\widehat{\lambda}^{2}_{n}$ is more efficient and can achieve better statistical power than using the first estimator $\widehat{\lambda}^{1}_{n}$ with the standard $t$-test based on asymptotic theory. We introduce the concept of the Efficient Tail Hypothesis (ETH), an analogue of the Efficient Market Hypothesis, to investigate the directional symmetry behavior of the market in its tails.
@@
In our data application, we investigate the market-wide tail efficiency of China’s futures market. Our empirical results suggest that the Efficient Tail Hypothesis (ETH) for China’s futures market is rejected under a significance level of $\alpha^{*}=0.01$. A deeper analysis of those tail-inefficient assets shows that negative extreme depreciation is more influential than positive extreme events. Moreover, we find evidence to support the momentum effect in extremes, which suggests that the market tends to be under-active in extreme situations. To capitalize on our findings on momentum effect in extremes, we construct an artificial dynamic portfolio consisting of those exceptional assets. The strategy is back-tested on out-of-sample data and shows a significant overall excess return. Those results suggest that most one-step forward extremal dependencies are directionally tail symmetric, while significant deviations indicate potential trading opportunities. Therefore, the analysis of directional tail dependence is useful for risk management and for constructing trading strategies in the financial market.
@@
Several directions can be explored in future research. One avenue is to extend the concept of directional tail dependence to higher dimensions. Additionally, current causal inference methods in extremes can be expanded to consider both upper and lower tails, using our new directional tail dependence measure. Last but not least, further investigation into simulation methods for time series data with specific directional tail dependence structures could enable the generation of time series data with varying degrees of directional asymmetry in financial applications.
@@
Supplementary Material for “The Efficient Tail Hypothesis: An Extreme Value Perspective on Market Efficiency”
Introduction
This Supplementary Material provides additional details and results for the main paper. Section A contains the proofs. Section B describes the dataset used in the Application section. Section C provides additional simulation results for the scenario with smaller sample sizes and lower exceedance thresholds. Finally, Section D presents further results for the empirical application.
@@ Appendix A Proofs
See 3.1
@@
[…]
@@
See 3.2
@@
[…]
@@
See 3.3
@@
for $\alpha\in[0,\pi/2],$ as shown in Illustration 6. Then, we express $\theta_{X^{+},Y^{-}}$ in terms of $\alpha$:
@@
See 3.4 We first show that $\lambda(X,Y)\in[-1,1]$, where $\lambda(X,Y)=-1$ and $\lambda(X,Y)=1$ are achieved when $\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y<-a]=1$ and $\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y>a]=1$, respectively. By Proposition 3.5, we have that $\lambda(X,Y)=2\{\sigma(X^{+},Y^{+})-\sigma(X^{+},Y^{-})\}$. As $\sigma(X^{+},Y^{+})\in[0,{1\over 2}]$ (and similarly for $\sigma(X^{+},Y^{-})$), we have that $\lambda(X,Y)\in[-1,1]$. The boundary case $\lambda(X,Y)=1$ implies $\sigma(X^{+},Y^{+})=0.5$ and $\theta_{X^{+},Y^{+}}=0$ which further implies $\sigma(X^{+},Y^{-})=0$ as $\theta_{X^{+},Y^{-}}\in[\frac{\pi}{2}-\theta_{X^{+},Y^{+}},\frac{\pi}{2}]$ by Proposition 3.3. The case $\sigma(X^{+},Y^{+})=0.5$ is achieved when
@@
See 3.5
@@
×$\times\frac{n}{\sigma_{n}\sqrt{k}}\left(\mathbb{E}[\omega^{+}_{x}\omega_{y}\mathbbm{1}\{a^{-1}(n/k)R\geq t^{-\frac{1}{2}}]-\frac{\lambda}{3}\mathbb{P}[a^{-1}(n/k)R\geq t^{-\frac{1}{2}}]\right)$

### v3

@@ 5.1 Overview and basic data analysis
We apply the framework developed in this paper to analyze the market-wide extremal dependence of China’s futures market. The data are represented as a multivariate time series of percentage price changes of futures for commodities with high liquidity. The return is calculated at 30-second intervals and the detailed procedure of data generation and access are given in the Appendix B. We consider a subset of the dataset containing 55 assets, which covers a wide range of commodities, including agricultural products, base metals, and ferrous metals. We select a one-year period of data from 2022-08-01 to 2023-07-31, which contains $T=\;$$110,985$ observations.
For each of the $55$ assets, we study bulk and tail temporal dependence by estimating their autocorrelation and extremogram (Davis et al.,, 2012) functions, respectively. The estimated (extremal) dependence decays rapidly, dropping below values of $0.05$ and $0.1$ (for the auto-correlation and extremogram, respectively) after $10$ time lags. Therefore, to remove temporal dependence from the data, we subsample every tenth observation. A detailed description of this dependence analysis is deferred to the Appendix D.1. After sub-sampling, we transform the data to be balanced regularly varying. In the Appendix D.2, we investigate the marginal tail behaviour of the data and find reasonable evidence to suggest that the upper and lower tails of each asset are regularly varying; we thus employ the first marginal transformation as described in Section 3.2. Unreported experiments found this choice to have negligible effect on the final results. We denote the transformed data by $\bm{Z}_{t}=(Z_{1,t},\dots,Z_{55,t})^{\top}$ for $t=1,\dots,T$.
The rest of this section is organized as follows: In Section 5.2, we visualize extremal dependence in the market using the extremal ball. In Section 5.3, we test the ETH for China’s futures market and interpret our findings. In Section 5.4, we capitalize on our test results to construct an artificial dynamic portfolio consisting of tail-inefficient assets and backtest an investment strategy on out-of-sample data. We also study the uncertainty arising from the marginal transformation via a bootstrap experiment, which is deferred to the Appendix D.4.
@@ 5.2 Visualization of market-wide extremal dependence
We observe that most of the points concentrate on the tail efficient line, covering a wide range of strengths of extremal dependence. This suggests that most one-step forward extremal dependencies are directionally tail symmetric. Although some extreme events appear to be closely linked to subsequent large movements in asset prices, this connection does not necessarily yield a clear directional signal. At best, it serves as an indicator of heightened volatility, implying that the market remains relatively efficient even under extreme conditions. A similar visualization is conducted for the contemporaneous extremal dependence structure which reveals the pattern of concurrence of extremal events across assets; we defer this to the Appendix D.3 for brevity.
Figure 4: Extremal ball for one-step forward dependencies in China’s futures market. Significant tail-inefficient pairs (p-values < 0.01) are circled based on the permutation test using $\hat{\lambda}^{2}_{n}$. Red: $(Z_{i,t},Z_{j,t+1})^{\top}$; Blue: $(-Z_{i,t},Z_{j,t+1})^{\top}$.
5.3 Testing the ETH for China’s futures market
@@
By analyzing whether $X_{i}$ and $Y_{j}$ belong to the same asset, we classify the significant pairs into two types: in-product pairs (i.e., $X_{i}$ and $Y_{j}$ belong to the same asset) and cross-product pairs (i.e., they belong to different assets). For the in-product significant pairs, all of them exhibit contrarian effects, i.e., ${\rm Sgn}(X_{i})\neq{\rm Sgn}(\lambda(X_{i},Y_{j}))$. Contrarian (momentum) effects are well-known phenomena in financial markets, reflecting a tendency for price changes to follow opposite (same) directional movements relative to recent returns of the same asset. The phenomenon of contrarian in extremes observed here may suggest that the market tends to overreact during extreme events and subsequently reverses in the following periods. For the cross-product pairs, the sign of the explanatory variable $X_{i}$ and the sign of $\lambda(X_{i},Y_{j})$ tend to be the same, which we refer to as co-directional lead-lag effects across different assets.
Statistics on all 60 significant pairs exhibit similar patterns: 41 pairs are driven by negative extreme events; There are only three in-product significant pairs: egg futures (driven by an extremal loss), natural rubber (driven by an extremal gain), and stainless steel (driven by an extremal gain). These three pairs are all contrarian in extremes. And for the remaining 57 cross-product pairs, 41 out of them are co-directional lead-lag effects across different assets.
Overall, the findings suggest that negative extreme movements are more influential than positive ones; contrarian effects in extreme dominates in in-product pairs; and co-directional lead-lag effects dominates in cross-product pairs.
Such statistics show that, in the context of extremes of China’s futures market, an asset’s own price often exhibits an initial overreaction, followed by a subsequent correction, which may form the basis of contrarian effects in extremes for tail-inefficient in-product pairs. As for the co-directional lead-lag effects in extremes for the cross-product pairs, it can be explained by the “slow information diffusion hypothesis” (Hou,, 2007). One plausible explanation is that certain assets swiftly incorporate relevant news, leading to initial price adjustments in explanatory assets, while others reflect the same information only after a delay. Moreover, the significant cross-product pairs may be interpreted as potential channels of tail risk transmission, thus connecting our analysis to the literature on spillover effects and lead-lad detection under market stress.
@@ 5.4 Exploiting the Efficient Tail Hypothesis analysis
For each pair satisfying $|\hat{\lambda}_{n}^{2}(X_{i},Y_{j})|>0$, we suppose that we enter into the market for asset $Y_{j}$ (long position if $\hat{\lambda}_{n}^{2}(X_{i},Y_{j})>0$, otherwise short position) when we observe an extreme value for asset $X_{i}$ (given it exceeds its 99.5% historical quantile for positive extremes and falls below its 0.5% quantile when considering negative extremes). We hold the asset for one period and then exit the market by closing the position. With the trading cost ignored, we backtest this strategy on the out-of-sample data from 2023-06-01 to 2023-07-31. The cumulative Profit and Loss for each of the 60 significant pairs is given in Figure 5, alongside the overall Profit and Loss for the portfolio. We also study the case when set significance level to $0.05$ under which we find 294 significant pairs. Both results show a positive Profit and Loss throughout, suggesting that a trading strategy based on tail inefficiencies has the potential to make excess returns relative to a baseline buy-and-hold portfolio consisting of the same 55 assets held with equal weights. This result is consistent with the bootstrap experiment, which is deferred to the Appendix D.4.
As stated by the Efficient Market Hypothesis, past information cannot be used to predict future price movements. However, our backtesting on out-of-sample data suggests that our trading strategy using tail-inefficient assets is profitable (under the ideal situation where the trading cost is not considered), which gives evidence against the Efficient Market Hypothesis. Therefore, the Efficient Tail Hypothesis, which is testable, can be viewed as an tail analogue of the Efficient Market Hypothesis with a stronger focus on extreme events.
Figure 5: Profit and Loss of a dynamic portfolio consisting of the 60 or 294 tail-inefficient pairs under significance level $\alpha^{*}=0.01$ (left) or $\alpha^{*}=0.05$ (right), respectively. Upper panels show the Profit and Loss for each pair and the lower panels show the running cumulative Profit and Loss for the portfolio. The $y$-axis is the cumulative Profit and Loss (%) and the $x$-axis is the time index for the two-month test period. The dashed blue curve represents the Profit and Loss of an equally-weighted portfolio of the 55 assets.
@@ 6 Conclusion
In this paper, we studied the pairwise directional tail dependence asymmetry for components of a random vector, $\bm{X}$. In comparison with traditional tail asymmetry, this new type of asymmetry focuses on describing asymmetry of tail dependence in two adjacent quadrants. Under balanced regular variation of $\bm{X}$ over the entire $\mathbb{R}^{p}$, we study the theoretical properties characterizing the dependence structure of $\bm{X}$ in all lower/upper tail quadrants, and we then create a novel tool, the extremal ball, to visualize the pairwise directional asymmetry of tail dependence within high-dimensional random vectors. Furthermore, we proved constraints on the tail dependence in two adjacent quadrants. Such a theoretical result suggests that knowledge of extremal dependence can aid in understanding the extremal dependence in adjacent quadrants of $\mathbb{R}^{p}$.
@@
We proposed a new measure, the directional tail dependence $\lambda(X,Y)$, to quantify the extent of directional tail symmetry. Two estimators are proposed for $\lambda(X,Y)$. We proved asymptotic normality for the first estimator; deriving the asymptotic properties for the second estimator remains future work. We further developed a permutation test for the null hypothesis $\mathcal{H}_{0}:\lambda(X,Y)=0$ and verified its effectiveness and power through simulation studies. We showcased that the permutation test with the second estimator $\widehat{\lambda}^{2}_{n}$ is more efficient and can achieve better statistical power than using the first estimator $\widehat{\lambda}^{1}_{n}$ with the standard $t$-test based on asymptotic theory. We introduce the concept of the Efficient Tail Hypothesis (ETH), an testable analogue of the Efficient Market Hypothesis, to investigate the directional symmetry behavior of the market in its tails. The statistical essence of ETH is directional tail symmetry between explanatory variables and future asset returns.
@@
In our data application, we investigate the market-wide tail efficiency of China’s futures market. Our empirical results suggest that the Efficient Tail Hypothesis (ETH) for China’s futures market is rejected under a significance level of $\alpha^{*}=0.01$. A deeper analysis of those tail-inefficient assets shows that negative extreme depreciation is more influential than positive extreme events. Moreover, we find evidence of co-directional lead-lag effects across different assets in extremes, which suggests that the market tends to be under-active in extreme situations. To capitalize on our findings, we construct an artificial dynamic portfolio consisting of those exceptional assets. The strategy is back-tested on out-of-sample data and shows a significant overall excess return. Those results suggest that most one-step forward extremal dependencies are directionally tail symmetric, while significant deviations indicate potential trading opportunities. Therefore, the analysis of directional tail dependence is useful for risk management and for constructing trading strategies in the financial market.
@@
Several directions can be explored in future research. One avenue is to extend the concept of directional tail dependence to higher dimensions. Additionally, current causal inference methods in extremes can be expanded to consider both upper and lower tails, using our new directional tail dependence measure. It is also of interest to study directional tail dependence during high-volatility periods, such as financial crises or at market opening and closing times. Last but not least, further investigation into simulation methods for time series data with specific directional tail dependence structures could enable the generation of time series data with varying degrees of directional asymmetry in financial applications.
Acknowledgments
We thank the Editor, the Associate Editor, and the two anonymous referees for their insightful comments, which have significantly improved the paper.
@@ Appendix A Proofs
Proposition 3.1. Let $\bm{X}^{*}=(X^{*}_{1},\dots,X^{*}_{p})^{\top}\in{\rm RV}^{p}(\alpha)$ with limit measure $v_{\bm{X}^{*}}(\cdot)$ satisfying $v_{\bm{X}^{*}}\left(\{\bm{x}\in\mathbb{E}^{p}:x_{i}>1\}\right)=c_{i}^{+}>0$ and $v_{\bm{X}^{*}}\left(\{\bm{x}\in\mathbb{E}^{p}:x_{i}<-1\}\right)=c_{i}^{-}>0$, for ${i=1,\dots,p}$. Then $\bm{X}=(X_{1},\dots,X_{p})^{\top}$, with $X_{i}=\max\{(c_{i}^{+})^{-1/\alpha}X^{*}_{i},0\}+\min\{(c_{i}^{-})^{-1/\alpha}X_{i}^{*},0\}$ for all ${i=1,\dots,p}$, satisfies $\bm{X}\in{\rm BRV}^{p}(\alpha)$.
@@
Condition i) follows by Lemma 6.1 of Resnick, (2007), which is a criterion for vague convergence to measures; It states that to show $\lim_{n\to\infty}n\mathbb{P}\{b_{n}^{-1}|\bm{X}|\in\cdot\}=\nu_{|\bm{X}|}(\cdot)$ on ${\mathbb{E}^{p}_{+}=\mathbb{R}_{+}\setminus\{\bm{0}\}}$, it is enough to show that $\lim_{n\to\infty}n\mathbb{P}\{b_{n}^{-1}|\bm{X}|\in[\bm{0},\bm{z}]^{c}\}=\nu_{|\bm{X}|}([\bm{0},\bm{z}]^{c})$ for all $\bm{z}:=(z_{1},\dots,z_{p})^{\top}\in[{0},\infty)^{p}\setminus\{\bm{0}\}$. We show that $\nu_{|\bm{X}|}(\cdot)$ can be expressed in terms of $\nu_{\bm{X}^{*}}(\cdot),$ which is the limit measure for $\bm{X}^{*}$. Consider
@@
Proposition 3.2 (Properties of $\bar{\bm{X}}$). Let $\bm{X}\in{\rm BRV}^{p}(2)$ and let $\bar{\bm{X}}:=\left((\bm{X}^{+})^{\top},(\bm{X}^{-})^{\top}\right)^{\top}$. Then $\bar{\bm{X}}\in{\rm RV}_{+}^{2p}(2)$ with limit measure $\nu_{\bar{\bm{X}}}(\cdot)$ that satisfies $\nu_{\bar{\bm{X}}}(\{\bm{x}\in\mathbb{E}^{2p}_{+}:x_{i}>1\})=1$ for $i=1,\dots,2p$, angular measure $H_{\bar{\bm{X}}}(\cdot)$ satisfying ${H_{\bar{\bm{X}}}(\mathbb{S}^{2p-1}_{+})}=2p$, and with $\sigma(X_{i}^{+},X_{i}^{-})=0$ for $i=1,\dots,p$ (orthogonality property).
@@
We first show that $\bar{\bm{X}}:=(\bar{X}_{1},\dots,\bar{X}_{2p})^{\top}\in{\rm RV}^{2p}_{+}(2)$ with limit measure $\nu_{\bar{\bm{X}}}(\cdot)$, which we express in terms of $\nu_{{\bm{X}}}(\cdot)$. By Lemma 6.1 of Resnick, (2007), we have, for all $\bm{\bar{z}}:=((\bm{z^{+}})^{\top};(\bm{z^{-}})^{\top})^{\top}\in[0,\infty)^{2p}\setminus\{\bm{0}\}$, that
@@
Proposition 3.3 (Extremal ball constraint). For $(X,Y)^{\top}\in{\rm BRV}^{2}(2)$ with $X^{+},Y^{+},Y^{-}\in\mathcal{V}^{q}_{+}$, the angles $\theta_{X^{+},Y^{+}}$ and $\theta_{X^{+},Y^{-}}$ satisfy the constraint $\frac{\pi}{2}-\theta_{X^{+},Y^{+}}\leq\theta_{X^{+},Y^{-}}\leq\frac{\pi}{2}$.
@@
for $\alpha\in[0,\pi/2],$ as shown in Figure 6. Then, we express $\theta_{X^{+},Y^{-}}$ in terms of $\alpha$:
@@
Proposition 3.4 (Properties of directional tail dependence $\lambda$). The directional tail dependence $\lambda(X,Y)$ satisfies ${\lambda(X,Y)\in[-1,1]}$, where ${\lambda(X,Y)=-1}$ and ${\lambda(X,Y)=1}$ are achieved when ${\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y<-a]=1}$ and ${\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y>a]=1}$, respectively. Furthermore, $\lambda(X,Y)$ is an odd function of $Y$, i.e., $\lambda(X,-Y)=-\lambda(X,Y)$ for all $Y$ (this property does not hold for $X$).
Proof of Proposition 3.4.
We first show that $\lambda(X,Y)\in[-1,1]$, where $\lambda(X,Y)=-1$ and $\lambda(X,Y)=1$ are achieved when $\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y<-a]=1$ and $\lim_{a\rightarrow\infty}\mathbb{P}[X>a\mid Y>a]=1$, respectively. By Proposition 3.5, we have that $\lambda(X,Y)=2\{\sigma(X^{+},Y^{+})-\sigma(X^{+},Y^{-})\}$. As $\sigma(X^{+},Y^{+})\in[0,{1\over 2}]$ (and similarly for $\sigma(X^{+},Y^{-})$), we have that $\lambda(X,Y)\in[-1,1]$. The boundary case $\lambda(X,Y)=1$ implies $\sigma(X^{+},Y^{+})=0.5$ and $\theta_{X^{+},Y^{+}}=0$ which further implies $\sigma(X^{+},Y^{-})=0$ as $\theta_{X^{+},Y^{-}}\in[\frac{\pi}{2}-\theta_{X^{+},Y^{+}},\frac{\pi}{2}]$ by Proposition 3.3. The case $\sigma(X^{+},Y^{+})=0.5$ is achieved when
@@
Proposition 3.5.   Consider the directional tail dependence $\lambda(X,Y)$ as in Definition 3.4. Then
@@
The Proof of Proposition 3.6 follows the same strategy as in Larsson and Resnick, (2012). The major difference comes from the estimator’s definition of an “exceedance”. In our case, we use $\mathbbm{1}\{r_{i}\geq r_{0}\}$ where $r_{0}=a(n/k)$ (same setting as used in Resnick, (2004)), while in Larsson and Resnick, (2012), they use $\mathbbm{1}\{r_{i}\geq r_{(k)}\}$ where $r_{(k)}$ is the $k$-th order statistic of $\left\{r_{i}\right\}_{i=1}^{n}$.
Proposition 3.6 (Asymptotic normality of $\hat{\lambda}^{1}_{n}$). Let $(X,Y)^{\top}\in{\rm BRV}^{2}(2)$ with ${\omega^{+}_{x}=\frac{X^{+}}{R}}$, ${\omega_{y}=\frac{Y}{R}}$, where $X^{+}=\max\{X,0\}$ and $R=||(X^{+},Y)^{\top}||_{2}$. Let $a(n)$ be a function satisfying $a(n)\rightarrow\infty$ and $n\mathbb{P}\{R>a(n)\}\rightarrow 1$ as $n\to\infty$. Define the “bias” process by
@@
×$\times\frac{n}{\sigma_{n}\sqrt{k}}\left(\mathbb{E}[\omega^{+}_{x}\omega_{y}\mathbbm{1}\{a^{-1}(n/k)R\geq t^{-\frac{1}{2}}\}]-\frac{\lambda}{3}\mathbb{P}[a^{-1}(n/k)R\geq t^{-\frac{1}{2}}]\right)$